solve the system of questions.


1/x-1/y+1/z=4


2/x+2/y+1/z=2


3/x+1/y-2/z=-2

Grade:

1 Answers

Chetan Mandayam Nayakar
12 years ago

let (1/x)=u, (1/y)=v and (1/z)=w, solving the system of equations in u,v,w and finding their values is ridiculously easy.